What is the best method to solve the system of equations?
2x + 3y = 8
4x - 6y = 2
I think that the best method would elimination with addition or multiplication. The other options are substitution and graphing.

Found 2 solutions by fractalier, MathTherapy:
Answer by fractalier(6550)

In this case, elimination is the best (not always, however)...
Here you would multiply the top one by 2 and add it to the bottom one...

Answer by MathTherapy(10557)


What is the best method to solve the system of equations?
2x + 3y = 8
4x - 6y = 2
I think that the best method would elimination with addition or multiplication. The other options are substitution and graphing.
Best method is to DIVIDE eq (ii) by 2. This will give you: 2x - 3y = 1 ----- eq (iii)
Now you can ADD eqs (iii) & (i) to find x. Then use value of x in any of the 2 original equations to find value of y